I have a test question that reads the question says:
SOLVE THE EQUATION
1.6x(squared)-24=0
and next time I have an answer may I put more than one question? Found 2 solutions by richwmiller, solver91311:Answer by richwmiller(17219) 
 It is best to only have one problem per question. There are many reasons.
Do not shout! Using capital letters is shouting.
What is squared? just the 1.6 x^2 or (1.6x)^2
for now 1.6x^2-24=0
divide by 8
.2x^2-3=0
.2x^2=3
divide by .2
x^2=3/.2
x^2=30/2
x^2=15
x=+\-sqrt(15)
which will be just a little less than 4 since 4^2=16